OpenStudy (anonymous):

Caleb deposited money into a savings account each week. The table shows the balance in his account after different numbers of weeks. Caleb graphed this information. What is the slope of the line that represents this situation?

OpenStudy (anonymous):

Number of weeks 3 6 9 15 Account balance($) 135 270 405 675

OpenStudy (anonymous):

slope = change in balance / change in weeks = (675 - 135)/(15 - 3) = 540/12 = 45

OpenStudy (cggurumanjunath):